OpenStudy (gina):

{xsin5xdx ..solve the intergratio

OpenStudy (amistre64):

your missing a -5; so id take one form -5/-5

OpenStudy (jamesj):

Use integration by parts. Let u = x and v' = sin(5x). Can you see what to do now?

OpenStudy (amistre64):

.... i dint think it was out of order lol

OpenStudy (amistre64):

sin(5x) comes from at least cos(5x) Dx (cos(5x)) = -5 sin(5x) ... ack, im missing an x from that . DOH!!

OpenStudy (amistre64):

good thing i think it through ... bad thing i think it through afterwards ;)

OpenStudy (gina):

@ james can u try to continue after that kuz ima finding it hard

OpenStudy (jamesj):

\[\int\limits u v' = u v - \int\limits u'v\] There must be an example in your notes or your text book of using this rule. Have a look at that example, line by line and then see how to apply it here. I've given you already what the u and v would be.

OpenStudy (gina):

@ james thanx so much ,cud also help me some examples on double intergrations??

OpenStudy (gina):

@ robtobey nice answer too but some steps are missing